If a o b is defined to be 2a-3b and 5 o x = 1, then x = ?
The answer is 3. Why?

Answer by Edwin McCravy(20054)

If a o b is defined to be 2a-3b and 5 o x = 1, then x = ?
The answer is 3. Why?

a o b = 2*a - 3*b

Substitute 5 for a and x for b

a o b = 2*a - 3*b
5 o x = 2*5 - 3*x

We are told that this equals 1, so

5 o x = 2*5 - 3*x =  1
          10 - 3x =  1
              -3x = -9
                x =  3

Checking:

5 o x = 2*5 - 3*x
5 o 3 = 2*5 - 3*3 = 10 - 9 = 1

Edwin